OpenStudy (highschoolmom2010):

A basketball player makes 80% of her free throws attempts. Find the average number of free throw attempts needed in order to make 3 free throws in a row.

OpenStudy (anonymous):

i think you can solve this via \(.8\times x=3\)

OpenStudy (anonymous):

My interpretation:\[ X \sim Binomial[n,p] = Binomial[n,0.80] \]We want \(n\) such that: \[ E[X] = np = n(0.8) = 3 \]

OpenStudy (anonymous):

However \(n\) has to be an integer, so you might only be able to get an average slightly less or over.

OpenStudy (highschoolmom2010):

im lost on this to be honest

OpenStudy (anonymous):

Do you know what the binomial probability distribution is?

OpenStudy (anonymous):

@wio didn't we write the same thing?

OpenStudy (anonymous):

\[.8x=3\iff x=\frac{8}{.3}\]

OpenStudy (anonymous):

I don't get your reasoning

OpenStudy (anonymous):

it is exactly what you wrote, cept i used \(x\) and you used \(n\) unless i misread your answer

OpenStudy (ybarrap):

Yes, both of yours are identical, it's the mean of the binomial random variable

OpenStudy (tkhunny):

Notice how the three made free throws are not necessarily consecutive.

OpenStudy (highschoolmom2010):

yep im still lost

OpenStudy (ybarrap):

@highschoolmom2010 you are given that the chance of a successful free throw is 80%. This means that the chance of failing is 20%. There are two probability distributions you should be considering: Binomial and Geometric. \(\underline{Binomial}\) When you have a random process where there are only 2 options, success or failure, we can use what is called the "binomial distribution". This same distribution can predict the number of heads in 5 tosses (for example), because in each attempt there are only two options, heads or tails. So whenever you encounter a random process that can only take on 2 values, think binomial. For example, if you wanted to know chance of making only 1 free throws in three tries, this would be \(0.80\times0.20\times0.20\) and since the free throw could have been made in either the 1st, 2nd or third attempt, we multiply this number times 3. If you can model a process by the binomial distribution, calculating the average number of successes in \(n\) tries is extremely easy. It is just the probability of success times the number of attempts: \(0.85\times n\) \(\underline{Geometric}\) There is another distribution type that take on two values, success and failure. That is the geometric distribution. This is the one where you want to find the number of \(consecutive\) unsuccessful attempts before you get a single successful one. The thing that should trigger you to use the geometric distribution is something like, "consecutive" or "in a row." The binomial models the chance of getting 3 free throws out of say 5, but not necessarily "in order." Your problem seems to be one of the geometric type. But we need to modify the parameters a little to make it "fit" the geometric. Take the tossing of a coin. The probability, we call it \(p\), of a head is 1/2. So the average number of tosses you need to get a head is \(\cfrac{1}{p}\). So, we need to toss the coin 2 times, \(on~average\) before we get a head. So for your problem, we need to find p. We will call a success "three consecutive free throws" and a failure otherwise. The chance of 3 consecutive free throws is \(0.80\times0.80\times0.80=0.80^3\). So, \(p=0.80^3\). And the average number of throws before we get three free throws in a row is \(\cfrac{1}{p}=\cfrac{1}{0.80^3}\approx1.95\). Each of these represents 3 free throw attempts, so we multiply by 3 to get the average number of free throw attempts to get 3 in a row: \(\cfrac{3}{0.80^3}\approx6\). For the 1st 2 throws, there is 0 chance of 3 in a row. So, when we computed 1.95 above, it is obvious that that could not be our final answer. It is not until the 3rd free-throw attempt, that "success" is even possible. So, chance of "success" starts on the 3rd free throw attempt, but on average we need 6 tries to get 3 successful free throws \(in~ a~ row\). If it were certain that we would make a free throw, then p=1 and 1/p = 1 is the average number of throws needed to get three in a row. But then we need to multiply by 3 to get the actual number of attempts. You can experiment with different values of p to see if the averages you get makes sense. For example, if each free throw had a 50% chance of success then you would need 24 attempts to get 3 in a row, on average. But with p = 90%, then you would need only between 4 and 5 attempts, on average, which is less than 6 (for your problem), where p= 80%. I hope this helps.

OpenStudy (highschoolmom2010):

@ybarrap yes that helpes

OpenStudy (ybarrap):

One last thing that I thought would be interesting to note. If you compare the number of attempts needed to get 3 successful free throws, regardless of order, we can use the binomial and that would be about 4 attempts, on average (because \(n\times.80=3\implies n=3.2\)) This makes sense since you would expect that it would be easier (and thus less attempts required) to make 3 free throws in any order, versus 3 free throws consecutively. This is like asking, which would take less time , to toss the coin until you get 3 heads in a row (geometric) or to toss the coin until you get 3 heads (binomial). Try it if you don't know -- toss a coin several times, chances are that you will get 3 heads in any order more frequently versus 3 heads in a row. Remember we are talking about averages, so you would have to try many times. So 4 (binomial average) < 6 (geometric average), which is what we would expect using this argument.